# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 1.5. Organization of the Paper
In Section 2, we introduce the Askey-Wilson process and the continuous dual Hahn process, and state results and definitions pertaining to the asymptotics of q-pochammer symbols. In Section 3, we state the main technical lemma, 3.1, that the Laplace transforms of the open ASEP height functions $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ converge pointwise to the Laplace transform of the open KPZ stationary measure when $u+v>0$ and $w,r>0$. The remaining parts of Section 3 demonstrate that the scaling limit of the Askey-Wilson measures produces the continuous dual Hahn process under the scaling assumptions in 1.1. In Section 4 we prove 3.1. In Section 5, we combine these results to prove 1.3.
@@ 1.6. Acknowledgements
The author thanks Ivan Corwin for suggesting the problem, and Ivan Corwin and Alisa Knizel for helpful discussions about their work. The author also thanks Włodek Bryc and Yizao Wang for helpful comments on a draft. The author’s research was supported by Ivan Corwin’s NSF grant DMS-1811143 and the Fernholz Foundation’s “Summer Minerva Fellows” program
@@ 2. Background
In Section 2.2 and Section 2.3 we will characterize the Askey-Wilson process and the continuous dual Hahn process, which are used to describe the stationary measures of the open ASEP and open KPZ, respectively. In Section 2.4 we will quote asymptotic results on q-pochammer symbols, proved in [CK24], which will be necessary for our argument. In Section 2.5, we formally state the result that obtains weak convergence of random variables from pointwise convergence of their Laplace transforms.
@@ 2.2. The Askey-Wilson Process
The stationary measure for the open ASEP can be described in terms of a stochastic process called the Askey-Wilson process. This process first appeared in the work of Bryc and Wesolowski [BW17], who formulated the probability generating function of the open ASEP in terms of the Askey-Wilson process $\mathbb{Y}_{t}$, see 2.1,
@@
In this paper, we will use the Askey-Wilson formula to characterize the Laplace transform of the open ASEP height function. We recall the definition of the diffusively scaled height function, $h^{(N)}(x)$ from 1.1. We can write its multipoint Laplace transform as
(2.1) $\phi^{(N)}_{u,v,w,r}(\textbf{c},\textbf{x})$ $=\left\langle e^{-\sum_{k=1}^{d}c_{k}\left(h^{(N)}_{u,v,w,r}(x_{k}^{(N)})-h^{(N)}_{u,v,w,r}(0)\right)}\right\rangle_{N}$
@@
(2.2) x $:=(x_{0},...,x_{d+1}),\text{ where }0=x_{0}<x_{1}<...<x_{d}\leq x_{d+1}=1,$
c $:=(c_{1},...,c_{d}),\text{ where }c_{1},...,c_{d}>0,$
@@
We will also use the notation $n_{k}=\lfloor Nx_{k}\rfloor$, $s_{k}^{(N)}=N^{-\frac{1}{2}}(s_{k}+c)$ and $x_{k}^{(N)}=N^{-1}\lfloor Nx_{k}\rfloor$. The rest of this subsection is devoted to defining the Askey-Wilson process.
Definition 2.1 (Section 1.3 [BW17]).
@@
(2.3) $\pi_{t}(V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right),$
$\pi_{s,t}(x,V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\sqrt{\frac{s}{t}}(x+\sqrt{x^{2}-1}),\sqrt{\frac{s}{t}}(x-\sqrt{x^{2}-1})\right).$
@@
$AW^{c}(x;a,b,c,d,q)$ $:=\mathbf{1}_{|x|<1}\frac{(q,ab,ac,ad,bc,bd,cd)_{\infty}}{2\pi(abcd)_{\infty}\sqrt{1-x^{2}}}\bigg{|}\frac{(e^{2i\theta_{x}})_{\infty}}{(ae^{i\theta_{x}},be^{i\theta_{x}},ce^{i\theta_{x}},de^{i\theta_{x}})_{\infty}}\bigg{|}^{2}$
@@
We refer to $AW^{c}(x;a,b,c,d,q)$ as the continuous density at $x$ and $AW^{d}(y;a,b,c,d,q)$ as the discrete mass at $y$. When we allow the parameter $q$ to scale in $N$, we will denote the corresponding one-point and transition densities by $\pi_{s}^{(N)}(\cdot)$ and $\pi_{s,t}^{(N)}(\cdot,\cdot)$.
@@
(2.4) $AW^{d}(y_{0}^{a};a,b,c,d,q)$ $:=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}},$
@@
2.2.1. One Point Distribution Atoms
In 2.1, the one-point distribution $\pi_{t}(V)$ takes the form of an Askey-Wilson measure $AW(V;a,b,c,d,q)$ with parameters
@@ 2.2.1. One Point Distribution Atoms
In the work of Corwin and Knizel [CK24], which we recover by setting $w=r=1$, the domain in $t$ on which convergence occurs is restricted to an interval $(-2,2)$ in order to prevent atoms from arising from cases $(2)$ and $(4)$. When $w$ and $r$ are allowed to vary, there is a comparable domain in which there are no atoms arising from cases $(2)$ and $(4)$. For any $t\in(-2w,2r)$ we will avoid all such atoms. In particular, since $w,r>0$, we note that $w+r>0$, and therefore $(-2w,2r)$ is an open interval. As a reminder for why this is necessary, note that establishing convergence in distribution through the Laplace transform, as described in 2.6, requires convergence on an open set.
@@ 2.2.2. Atoms and Time Reversal
In the discussion of the asymptotics of the transition probabilities, we center and rescale the Askey-Wilson Process. If the original Askey-Wilson process is $\mathbb{Y}$, characterized by $q,A,B,C,D,$ then we define the tangent process $\hat{\mathbb{Y}}^{(N)}_{s}=2N(1-\mathbb{Y}_{q^{s}})$. This parameterization involves a time reversal of $\mathbb{Y}$, and the transition probabilities of this process are related to the transition probabilities of the original process by a conjugation, which we will describe below.
@@
We denote the measure associated to $\hat{\mathbb{Y}}_{s}$ by $\hat{\pi}_{s}^{(N)}$. As before, this measure has a discrete support and a continuous support, given by
$\hat{S}_{s}^{(N),c}=\hat{S}^{(N),c}:=[0,4N]$ $\hat{S}_{s}^{(N),d}:=\left\{y\in\mathbb{R}:1-\frac{y}{2N}\in S^{d}_{q^{s}}\right\}.$
@@
In the continuous component of the support, we will additionally emphasize the dependence on the parameters $w,v,u,r$ by writing $\hat{\pi}_{s;w,v,u,r}^{(N)}$. In this notation, we can define $\hat{\pi}_{t}$ and $\hat{\pi}_{s,t}$ in terms of $\pi_{t}$ and $\pi_{s,t}$ via the following transformations.
$\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $:=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right),$ $y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{t}^{(N),d}(y)$ $:=\pi_{q^{t}}^{(N),d}\left(1-\frac{y}{2N}\right),$ $y\in\hat{S}^{(N),d}_{t}.$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s;v,w,u,r}^{(N),c},y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),d},y\in\hat{S}_{t}^{(N),c},$
$\hat{\pi}_{s,t;v,w,u,r}^{(N),c,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),c},y\in\hat{S}_{t}^{(N),d},$
$\hat{\pi}_{s,t}^{(N),d,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\hat{S}_{s}^{(N),d},y\in\hat{S}_{t}^{(N),d}.$
@@
The discrete support of $\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{y}{2N},\cdot\right)$ is not equal to the set $\hat{S}_{s}^{(N)}$, and likewise for $\pi_{q^{t},q^{s}}\left(\cdot,1-\frac{x}{2N}\right)$. These transition probabilities are set to be zero when $x$ and $y$ are not located, respectively, in $\hat{S}_{s}^{(N)}$ or $\hat{S}_{t}^{(N)}$.
The tangent process $\hat{\mathbb{Y}}_{s}^{(N)}$ has a symmetry on the continuous support, given by
(2.5) $\hat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\hat{\pi}_{t;w,v,r,u}^{(N),c}(4N-y).$
@@
The Askey-Wilson has a limiting process called the continuous dual Hahn process, which was defined by Corwin and Knizel [CK24] in order to characterize the stationary measure of the open KPZ.
@@ 2.3. The Continuous Dual Hahn Process
In this section, we give the definition of the continuous dual Hahn process. The definition in this paper is restricted to the time interval $t\in[0,C_{u,r})$ where $C_{u}=2r$ if $u\leq 0$ or $u\geq 1$ and $C_{u}=\min{\{2u,2r\}}$ if $u\in(0,1)$. It is possible to define the continuous dual Hahn process for all $t\in\mathbb{R}$, but the more general formulas are significantly more complicated, and not necessary for the arguments in this paper. In definitions which follow, $\mathfrak{p}_{t}(\cdot)$ will not be a probability distribution, but instead a positive distribution of infinite mass, while $\mathfrak{p}_{s,t}(\cdot,\cdot)$ will be a probability distribution.
Definition 2.2 (Section 6.2 [CK24]).
The continuous dual Hahn process is the process with one-point and transition measures given by
@@
where $\mathfrak{p}^{c}_{t}(\cdot)$ is the one-point continuous dual Hahn density on the continuous support of the process, and $\mathfrak{p}^{d}_{t}(\cdot)$ is likewise the discrete mass defined on the discrete support of the process, and $\textsc{CDH}(V;a(x),b(x),c(x);\mathcal{F})$ is the continuous dual Hahn measure (See 2.3) for $x\in\mathcal{F}$ and $a(x),b(x),c(x)$ constants depending on $x$. These measures are supported on a set with a continuous component, $S^{c}_{t}$, and a discrete component, $S^{d}_{t}$, which is further broken up into two parts. We define these supports as
$S^{c}_{t}:=\mathbb{R}_{>0}$ $S^{d,u}_{t}:=\bigcup_{j=0}^{\lfloor-u+\frac{t}{2}\rfloor}\{x_{j}^{u}(t)\},$ $S_{t}^{d,v}:=\bigcup_{j=0}^{\lfloor-v-\frac{t}{2}\rfloor}\{x_{j}^{v}(t)\}.$
@@
(2.7) $\mathfrak{p}_{t}^{c}(x):=\frac{(u+v+1)(u+v)}{8\pi}\cdot\frac{|\Gamma\left(v+\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)\Gamma\left(u-\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)|^{2}}{\sqrt{x}|\Gamma(i\sqrt{x})|^{2}}\mathbf{1}_{x>0}.$
@@
(2.8) $\mathfrak{p}_{t}^{d}(x_{j}^{u}(t)):=\frac{\Gamma(v-u+t)\Gamma(v+u+2)}{\Gamma(-2u+t)}\cdot\frac{\left(u+j-\frac{t}{2}\right)[2u-t,v+u]_{j}}{\left(u-\frac{t}{2}\right)[1,1-v+u-t]_{j}},$
(2.9) $\mathfrak{p}_{t}^{d}(x_{j}^{v}(t)):=\frac{\Gamma(u-v-t)\Gamma(v+u+2)}{\Gamma(-2v-t)}\cdot\frac{\left(v+j+\frac{t}{2}\right)[2v+t,v+u]_{j}}{\left(v+\frac{t}{2}\right)[1,1-v+u-t]_{j}}.$
To define the transition densities, we use $\mathcal{S}_{t}$ to denote the union of the continuous and discrete supports at time $t$. In the definition which follows, we use the notation $\textsc{CDH}(V;a,b,c;\mathcal{F})$ to denote the continuous dual Hahn measure, which is introduced in 2.3. For $x\in S_{s}^{c}$, we define
(2.10) $\mathfrak{p}_{s,t}(x,V):=\textsc{CDH}\left(V;u-\frac{t}{2},\frac{t-s}{2}+i\frac{\sqrt{x}}{2},\frac{t-s}{2}-i\frac{\sqrt{x}}{2};\mathcal{S}_{t}\right).$
If $u-\frac{s}{2}<0$ (and consequently, $S_{s}^{d}=S_{s}^{d,u}$, then for $j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]$,
(2.11) $\mathfrak{p}_{s,t}(x_{j}^{u}(s),V):=\textsc{CDH}\left(V;u-\frac{t}{2},-u+\frac{t}{2}-j,u+\frac{t}{2}-s+j;\mathcal{S}_{t}\right).$
@@
(2.12) $\mathfrak{p}_{s,t}(x_{j}^{v}(s),V):=\textsc{CDH}\left(V;v+j+\frac{t}{2},\frac{t}{2}-s-v-j,u-\frac{t}{2};\mathcal{S}_{t}\right).$
In all other cases, we define $\mathfrak{p}_{s,t}$ to be zero.
Definition 2.3 (Definitions 6.3, 6.4 [CK24]).
Assume $a,b,c\in\mathbb{R}$ with $a+b,a+c>0$ or $a\in\mathbb{R}$ and $b=\overline{c}$ with $Re(b)=Re(c)>0$. For such $a,b,c$ we define a density function
@@
On all other arguments, $\textsc{CDH}^{d}(\cdot;a,b,c;\mathcal{F})$ is zero. Finally, we define the full continuous dual Hahn measure $\textsc{CDH}(V;a,b,c;\mathcal{F})$ in the following three cases:
@@
Finally, we note the following property of the continuous dual Hahn process.
Lemma 2.4 (Lemma 6.9 [CK24]).
For $u,v\in\mathbb{R}$ with $u+v>0$ and $0\leq s<t<w<C_{u,v}$ and for any Borel set $V\subset\mathbb{R}$,
$\int_{\mathbb{R}}\mathfrak{p}_{s}(dm)\mathfrak{p}_{s,t}(m,V)=\mathfrak{p}_{t}(V)$ $\int_{\mathbb{R}}\mathfrak{p}_{s,t}(m,dr)\mathfrak{p}_{t,w}(r,V)=\mathfrak{p}_{s,w}(m,V)$

### v2

@@ 1.5. Organization of the Paper
In Section 2, we introduce the Askey-Wilson process and the continuous dual Hahn process, and state results and definitions pertaining to the asymptotics of $q$-Pochhammer symbols. In Section 3, we state the main technical lemma, 3.1, that the Laplace transforms of the open ASEP height functions $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}_{N\in\mathbb{N}}$ converge pointwise to the Laplace transform of the open KPZ stationary measure when $u+v>0$ and $w,r>0$. The remaining parts of Section 3 demonstrate that the scaling limit of the Askey-Wilson measures produces the continuous dual Hahn process under the scaling assumptions in 1.1. In Section 4 we prove 3.1. In Section 5, we combine these results to prove 1.1.
@@ 1.6. Acknowledgements
The author thanks Ivan Corwin for suggesting the problem, and Ivan Corwin and Alisa Knizel for helpful discussions about their work. The author also thanks Włodek Bryc and Yizao Wang for helpful comments on a draft. The author’s research was supported by Ivan Corwin’s NSF grant DMS-1811143 and the Fernholz Foundation’s “Summer Minerva Fellows” program.
@@ 2. Background
In Section 2.2 and Section 2.3 we will characterize the Askey-Wilson process and the continuous dual Hahn process, which are used to describe the stationary measures of the open ASEP and open KPZ equation, respectively. In Section 2.4 we will quote asymptotic results on $q$-Pochhammer symbols, proved in [CK24], which will be necessary for our argument. In Section 2.5, we formally state the result that obtains weak convergence of random variables from pointwise convergence of their Laplace transforms.
@@ 2.2. The Askey-Wilson Process
The stationary measure for the open ASEP can be described in terms of a stochastic process called the Askey-Wilson process. This process first appeared in [BW17], which formulated the probability generating function of the open ASEP in terms of the Askey-Wilson process $\mathbb{Y}_{t}$, see 2.1,
@@
In this paper, we will use the Askey-Wilson formula to characterize the Laplace transform of the open ASEP height function. We recall the definition of the diffusively scaled height function, $h^{(N)}(x)$ from 1.1. We can write its multi-point Laplace transform as
$\phi^{(N)}_{u,v,w,r}(\vec{c},\vec{x})$ $=\left\langle e^{-\sum_{k=1}^{d}c_{k}\left(h^{(N)}_{u,v,w,r}(x_{k}^{(N)})-h^{(N)}_{u,v,w,r}(0)\right)}\right\rangle_{N}$ (2.1)
@@
$\vec{x}$ $:=(x_{0},...,x_{d+1}),\text{ where }0=x_{0}<x_{1}<...<x_{d}\leq x_{d+1}=1,$ (2.2)
$\vec{c}$ $:=(c_{1},...,c_{d}),\text{ where }c_{1},...,c_{d}>0,$
@@
We will also use the notation $n_{k}=\lfloor Nx_{k}\rfloor$, $s_{k}^{(N)}=N^{-\frac{1}{2}}s_{k}$ and $x_{k}^{(N)}=N^{-1}\lfloor Nx_{k}\rfloor$. The rest of this subsection is devoted to defining the Askey-Wilson process, following the definitions in [BW17].
Definition 2.1.
@@
$\pi_{t}(V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right),$ (2.3)
$\pi_{s,t}(x,V)$ $:=AW\left(V;A\sqrt{t},B\sqrt{t},\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right),\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)\right).$
@@
$AW^{c}(x;a,b,c,d,q)$ $:=\mathbf{1}_{|x|<1}\frac{(q,ab,ac,ad,bc,bd,cd)_{\infty}}{2\pi(abcd)_{\infty}\sqrt{1-x^{2}}}\bigg{|}\frac{(e^{2i\theta_{x}})_{\infty}}{(ae^{i\theta_{x}},be^{i\theta_{x}},ce^{i\theta_{x}},de^{i\theta_{x}})_{\infty}}\bigg{|}^{2},$
@@
We refer to $AW^{c}(x;a,b,c,d,q)$ as the continuous density at $x$ and $AW^{d}(y;a,b,c,d,q)$ as the discrete mass at $y$. For an open ASEP of lattice size $N$, with parameters scaled according to 1.1, we will denote the corresponding marginal and transition laws by $\pi_{s}^{(N)}(\cdot)$ and $\pi_{s,t}^{(N)}(\cdot,\cdot)$.
@@
$AW^{d}(y_{0}^{a};a,b,c,d,q)$ $:=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}},$ (2.4)
@@
2.2.1. Atoms in the Marginal Law
In 2.1, the marginal distribution $\pi_{t}(V)$ takes the form of an Askey-Wilson measure $AW(V;a,b,c,d,q)$ with parameters
@@ 2.2.1. Atoms in the Marginal Law
We recover the case of [CK24] by setting $w=r=1$. In this setting, the domain in $t$ on which convergence occurs is restricted to an interval $(-2,2)$ in order to prevent atoms associated to the parameters $b$ in case $(2)$ and to $d$ in case $(4)$. When $w$ and $r$ are allowed to vary, there is a comparable domain in which there are no atoms arising from cases $(2)$ and $(4)$. For any $t\in(-2w,2r)$ we will avoid all such atoms. In particular, since $w,r>0$, $(-2w,2r)$ is an open interval. As a reminder for why this is necessary, note that establishing convergence in distribution through the Laplace transform, as described in 2.4, requires convergence on an open set.
@@ 2.2.2. Atoms and Time Reversal
In the discussion of the asymptotics of the transition probabilities, we center and rescale the Askey-Wilson Process. If the original Askey-Wilson process is $\mathbb{Y}$, characterized by $q,A,B,C,D,$ then we define the tangent process
$\widehat{\mathbb{Y}}^{(N)}_{s}=2N(1-\mathbb{Y}_{q^{s}}).$
This parameterization involves a time reversal of $\mathbb{Y}$, and the transition probabilities of this process are related to the transition probabilities of the original process by a conjugation, which we will describe below. Though we will not take this approach, it is relevant to note that it is possible to avoid performing a time reversal by taking the time transformation $t\mapsto e^{2t/\sqrt{N}}$ instead of $t\mapsto q^{t}$. Other papers which employ the Askey-Wilson process, for instance [WY24], take this approach.
@@
We denote the measure associated to $\widehat{\mathbb{Y}}_{s}$ by $\widehat{\pi}_{s}^{(N)}$. As before, this measure has a discrete support and a continuous support, given by
$\widehat{S}_{s}^{(N),c}=\widehat{S}^{(N),c}:=[0,4N],$ $\widehat{S}_{s}^{(N),d}:=\left\{y\in\mathbb{R}:1-\frac{y}{2N}\in S^{d}_{q^{s}}\right\},$
@@
where $S^{d}_{q^{s}}$ is the collection of atoms in $S^{d}(q^{v+s/2},-q^{1+s/2},q^{u-s/2},q^{1-s/2},q)$, this set is described in Section 2.2. In the continuous component of the support, we will additionally emphasize the dependence on the parameters $w,v,u,r$ by writing $\widehat{\pi}_{s;w,v,u,r}^{(N)}$. In this notation, we can define $\widehat{\pi}_{t}$ and $\widehat{\pi}_{s,t}$ in terms of $\pi_{t}$ and $\pi_{s,t}$ via the following transformations,
$\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)$ $:=\frac{1}{2N}\pi_{q^{t}}^{(N),c}\left(1-\frac{y}{2N}\right),$ $y\in\widehat{S}_{t}^{(N),c},$
$\widehat{\pi}_{t}^{(N),d}(y)$ $:=\pi_{q^{t}}^{(N),d}\left(1-\frac{y}{2N}\right),$ $y\in\widehat{S}^{(N),d}_{t}.$
$\widehat{\pi}_{s,t;v,w,u,r}^{(N),c,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\widehat{S}_{s}^{(N),c},y\in\widehat{S}_{t}^{(N),c},$
$\widehat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(x,y)$ $:=\frac{1}{2N}\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),c}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\widehat{S}_{s}^{(N),d},y\in\widehat{S}_{t}^{(N),c},$
$\widehat{\pi}_{s,t;v,w,u,r}^{(N),c,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,c}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),c}(1-\frac{x}{2N})},$ $x\in\widehat{S}_{s}^{(N),c},y\in\widehat{S}_{t}^{(N),d},$
$\widehat{\pi}_{s,t}^{(N),d,d}(x,y)$ $:=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\frac{\pi_{q^{t}}^{(N),d}(1-\frac{y}{2N})}{\pi_{q^{s}}^{(N),d}(1-\frac{x}{2N})},$ $x\in\widehat{S}_{s}^{(N),d},y\in\widehat{S}_{t}^{(N),d}.$
@@
The discrete support of $\widehat{\pi}_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},\cdot\right)$ is not equal to the set $\widehat{S}_{s}^{(N),d}$, and likewise for $\widehat{\pi}_{q^{t},q^{s}}\left(\cdot,1-\frac{x}{2N}\right)$. These transition probabilities are set to be zero when $x$ and $y$ are not located, respectively, in $\widehat{S}_{s}^{(N),d}$ or $\widehat{S}_{t}^{(N),d}$.
On its continuous support, $\widehat{\pi}^{(N),c}_{t;v,w,u,r}(\cdot)$ has the following symmetry,
$\widehat{\pi}_{t;v,w,u,r}^{(N),c}(y)=\widehat{\pi}_{t;w,v,r,u}^{(N),c}(4N-y).$ (2.5)
@@
The Askey-Wilson has a limiting process called the continuous dual Hahn process, which was defined in [CK24] in order to characterize the stationary measure of the open KPZ equation.
@@ 2.3. The Continuous Dual Hahn Process
In this section, we give the definition of the continuous dual Hahn process. The definition in this paper is restricted to the time interval $t\in[0,C_{u,r})$ where $C_{u}=2r$ if $u\leq 0$ or $u\geq 1$ and $C_{u}=\min{\{2u,2r\}}$ if $u\in(0,1)$. It is possible to define the continuous dual Hahn process for all $t\in\mathbb{R}$, but the more general formulas are significantly more complicated, and not necessary for the arguments in this paper. The definition of the continuous dual Hahn process for $u,v>0$ also appears in [Bry12]. In definitions which follow, $\mathfrak{p}_{t}(\cdot)$ will not be the density function of a probability distribution, but instead of a positive distribution of infinite mass, while $\mathfrak{p}_{s,t}(x,\cdot)$ will correspond to a probability distribution. We follow the definition from [CK24, Section 6.2].
Definition 2.2.
The continuous dual Hahn process is the process with marginal and transition measures given by
@@
where $\mathfrak{p}^{c}_{t}(\cdot)$ is the marginal continuous dual Hahn law on the continuous support of the process, and $\mathfrak{p}^{d}_{t}(\cdot)$ is likewise the discrete mass defined on the discrete support of the process, and $\textsc{CDH}(V;a(x),b(x),c(x);\mathcal{F})$ is the continuous dual Hahn measure (See 2.3) for $x\in\mathcal{F}$ and $a(x),b(x),c(x)$ constants depending on $x$. These measures are supported on a set with a continuous component, $S^{c}_{t}$, and a discrete component, $S^{d}_{t}$, which is further broken up into two parts. We define these supports as
$S^{c}_{t}:=\mathbb{R}_{>0},$ $S^{d,u}_{t}:=\bigcup_{j=0}^{\lfloor-u+\frac{t}{2}\rfloor}\{x_{j}^{u}(t)\},$ $S_{t}^{d,v}:=\bigcup_{j=0}^{\lfloor-v-\frac{t}{2}\rfloor}\{x_{j}^{v}(t)\}.$
@@
$\mathfrak{p}_{t}^{c}(x):=\frac{(u+v+1)(u+v)}{8\pi}\cdot\frac{|\Gamma\left(v+\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)\Gamma\left(u-\frac{t}{2}+i\frac{\sqrt{x}}{2}\right)|^{2}}{\sqrt{x}|\Gamma(i\sqrt{x})|^{2}}\mathbf{1}_{x>0}.$ (2.7)
@@
$\mathfrak{p}_{t}^{d}(x_{j}^{u}(t)):=\frac{\Gamma(v-u+t)\Gamma(v+u+2)}{\Gamma(-2u+t)}\cdot\frac{\left(u+j-\frac{t}{2}\right)[2u-t,v+u]_{j}}{\left(u-\frac{t}{2}\right)[1,1-v+u-t]_{j}},$ (2.8)
$\mathfrak{p}_{t}^{d}(x_{j}^{v}(t)):=\frac{\Gamma(u-v-t)\Gamma(v+u+2)}{\Gamma(-2v-t)}\cdot\frac{\left(v+j+\frac{t}{2}\right)[2v+t,v+u]_{j}}{\left(v+\frac{t}{2}\right)[1,1-v+u-t]_{j}}.$ (2.9)
To define the transition densities, we use $\mathcal{S}_{t}:=S_{t}^{c}\cup S_{t}^{d,u}\cup S_{t}^{d,v}$ to denote the union of the continuous and discrete supports at time $t$. In the definition which follows, we use the notation $\textsc{CDH}(V;a,b,c;\mathcal{F})$ to denote the continuous dual Hahn measure, which is introduced in 2.3. For $x\in S_{s}^{c}$, we define
$\mathfrak{p}_{s,t}(x,V):=\textsc{CDH}\left(V;u-\frac{t}{2},\frac{t-s}{2}+i\frac{\sqrt{x}}{2},\frac{t-s}{2}-i\frac{\sqrt{x}}{2};\mathcal{S}_{t}\right).$ (2.10)
When $V$ is contained in the discrete support, we use the notation $\mathfrak{p}_{s,t}^{c,d}$, and likewise when $V$ is contained in the continuous support we write $\mathfrak{p}_{s,t}^{c,d}$. If $u-\frac{s}{2}<0$ (and consequently, $S_{s}^{d}=S_{s}^{d,u}$), then for $j\in[[0,\lfloor-u+\frac{s}{2}\rfloor]]$,
$\mathfrak{p}_{s,t}(x_{j}^{u}(s),V):=\textsc{CDH}\left(V;u-\frac{t}{2},-u+\frac{t}{2}-j,u+\frac{t}{2}-s+j;\mathcal{S}_{t}\right).$ (2.11)
@@
$\mathfrak{p}_{s,t}(x_{j}^{v}(s),V):=\textsc{CDH}\left(V;v+j+\frac{t}{2},\frac{t}{2}-s-v-j,u-\frac{t}{2};\mathcal{S}_{t}\right).$ (2.12)
In both of the preceding cases, we use the notation $\mathfrak{p}_{s,t}^{d,d}$ and $\mathfrak{p}_{s,t}^{d,c}$ to denote the transition to the discrete and continuous supports, respectively. In all other cases, we define $\mathfrak{p}_{s,t}$ to be zero.
The next definition, of $\textsc{CDH}^{c}(x;a,b,c)$ and $\textsc{CDH}^{d}(x_{j};a,b,c;\mathcal{F})$ follows [CK24, Definitions 6.3 and 6.4].
Definition 2.3.
Assume $a,b,c\in\mathbb{R}$ with $a+b,a+c>0$ or $a\in\mathbb{R}$ and $b=\overline{c}$ with $Re(b)=Re(c)>0$. For such $a,b,c$ we define
@@
On all other arguments, $\textsc{CDH}^{d}(\cdot;a,b,c;\mathcal{F})$ is zero. Finally, we define the continuous dual Hahn measure $\textsc{CDH}(V;a,b,c;\mathcal{F})$ in the following three cases: